# Impact of existence and nonexistence of pivot on the coverage of empirical best linear prediction intervals for small areas

### v2

@@ 5.2 Further simulations on asymmetric cases
While some of our theoretical results for single bootstrap are based on the symmetry assumption, we also use simulation studies to assess our proposed parametric bootstrap method under asymmetric conditions. Specifically, for the same models and parameter choices as above, we change $\mathcal{G}$ to be a shifted exponential distribution (SE). Besides the seven prediction intervals mentioned before, we also include our proposed double parametric bootstrap intervals in this subsection and we expect they might improve the potential coverage error under asymmetric conditions. Below, the double bootstrap method based on the Fay-Herriot variance estimator is denoted by DB.FH, and the double bootstrap method based on the Prasad-Rao variance estimator is indicated by DB.PR. We keep bootstrap sample of size 400 in the first stage and apply two different sizes $B_{2}\in(50,100)$ in the second stage. Since for $B_{2}=50$ and $B_{2}=100$, we obtained very similar results. Moreover, the two patterns of $D_{i}$ gave us similar results. In the following, we provide detailed discussions only for $B_{2}=100$ under the pattern (i), but the results for $B_{2}=50$ are provided in the Appendix G.
@@
Table 4: Percentages of negative estimates of $A$ ($\hat{A}^{*}$)[$\hat{A}^{**}$] for different estimation methods and different patterns of $D_{i}$ under SE distribution.
@@ 6 Discussion
In this study, we put forward parametric bootstrap approaches to construct prediction intervals in the contexts of small area estimation under a general (non-normal) area-level model. Our simulation results show that the proposed single bootstrap method with the Fay-Herriot method of variance estimator performs well for all cases. Moreover, it is more efficient in terms of average lengths than the existing parametric bootstrap method proposed by Hall and Maiti (2006).
@@
When the number of area is small, there is a high likelihood to obtain negative estimates of the variance $A$ when applying Prasad-Rao method. Throughout our simulation studies, we notice that the likelihood of obtaining negative estimates of variance might affect the performance of the parametric bootstrap methods in terms of both coverage probabilities and average lengths. One might consider a better variance estimator, such as the Fay-Herriot estimator we used, or a sensible truncation of negative estimates. In this study, we arbitrarily truncate the negative estimates of the variance $A$ at 0.01 as similar to Datta et al. (2005). To this end, in the future we will consider an extension of adjusted maximum likelihood estimators of $A$ such as the ones considered by Li and Lahiri (2010) or Hirose and Lahiri (2018) to the model proposed in this paper.

### v3

@@ 5.2 Further simulations on asymmetric cases
While some of our theoretical results for single bootstrap are based on the symmetry assumption, we also use simulation studies to assess our proposed parametric bootstrap method under asymmetric conditions. Specifically, for the same models and parameter choices as above, we change $\mathcal{G}$ to be a shifted exponential distribution (SE). Besides the seven prediction intervals mentioned before, we also include our proposed double parametric bootstrap intervals in this subsection and we expect they might improve the potential coverage error under asymmetric conditions. Below, the double bootstrap method based on the Fay-Herriot variance estimator is denoted by DB.FH, and the double bootstrap method based on the Prasad-Rao variance estimator is indicated by DB.PR. We keep bootstrap sample of size 400 in the first stage and apply two different sizes $B_{2}\in(50,100)$ in the second stage. Since for $B_{2}=50$ and $B_{2}=100$, we obtained very similar results. Moreover, the two patterns of $D_{i}$ gave us similar results. In the following, we provide detailed discussions only for $B_{2}=100$ under the pattern (i), but the results for $B_{2}=50$ are provided in the Supplementary Material. The average computation time per simulation run for the proposed double bootstrap method is 22.41 seconds for $m=15,\;B_{2}=50$; 36.61 seconds for $m=15,\;B_{2}=100$; 57.60 seconds for $m=50,\;B_{2}=50$; and 97.10 seconds for $m=50,\;B_{2}=100$.
@@
Table 4: Percentages of negative estimates of AA ($\hat{A}^{*}$)[$\hat{A}^{**}$] for different estimation methods and different patterns of $D_{i}$ under SE distribution.
@@
6 Real Data Analysis
In this section, we use data from the 1989 Small Area Income and Poverty Estimates (SAIPE) program to demonstrate the proposed parametric bootstrap methods for constructing prediction intervals. The SAIPE program produces poverty statistics for various age groups at different levels of geographic aggregation for the 50 states and the District of Columbia. A general area-level model is applied to direct poverty estimates from the U.S. Current Population Survey (CPS), borrowing information from administrative records such as selected tax data from the IRS and participation data from the Supplemental Nutrition Assistance Program (SNAP, formerly known as the Food Stamp Program).
For the 5–17 age group, Bell and Huang (2006) detected that the 1989 CPS direct estimate for the state of Connecticut could be regarded as a potential outlier. To address potential outliers, they considered area-level models with $t$-distributed random effects. One such model is given by:
$y_{i}=\bm{x}_{i}^{\prime}\bm{\beta}+u_{i}+e_{i},\quad i=1,\ldots,51,$
where $u_{i}\overset{iid}{\sim}t_{\phi}(0,A)$ with 3 degrees of freedom (i.e., $\phi=3$) and $e_{i}\overset{ind}{\sim}N(0,D_{i})$. The SAIPE data provide the direct CPS estimated poverty rates for related children aged 5–17 ($y_{i}$), associated sampling variances ($D_{i}$), and four auxiliary variables ($\bm{x}_{i}$), including pseudo-poverty rates tabulated from IRS tax data, tax non-filer rates from IRS data, Food Stamp participation proportions, and previous census estimated poverty ratios.
By replacing the unknown parameters $\bm{\beta}$ and $A$ with the weighted least squares (WLS) and Fay-Herriot (FH) estimates, respectively, we can obtain the EBLUP, $\hat{\theta}_{i}$, of the poverty ratios for each state. Following Bell et al. (2007), we set $\alpha=90\%$. We then implement both the proposed single ($B_{1}=400$; t,SB) and double ($B_{2}=100$; t,DB) parametric bootstrap methods to obtain the 90 percent prediction intervals and compare them with those obtained the proposed single parametric bootstrap method under the case $k=\infty$ (i.e., $u_{i}$ follows a normal distribution; normal) and with direct confidence intervals (DIRECT). The computation times are 32.26 seconds and 136.63 seconds for the proposed single and double bootstrap methods, respectively.
Figure 1 presents the poverty ratio estimates and associated 90% prediction intervals for each state. For most states, the direct intervals are too wide to be informative. The lengths of the single bootstrap intervals based on normally distributed and $t$-distributed random effects are similar. The double bootstrap intervals are longer than the single bootstrap intervals and tend to contain them. This corresponds with the theory above that double bootstrap intervals could potentially offer better coverage than single bootstrap intervals.
Figure 1: 90% Prediction intervals and estimates for State 5-17 1989 poverty ratios.
@@ 7 Discussion
In this study, we put forward parametric bootstrap approaches to construct prediction intervals in the contexts of small area estimation under a general (non normal) area level model. Our simulation results show that the proposed single bootstrap method with the Fay-Herriot method of variance estimator performs well for all cases. Moreover, it is more efficient in terms of average lengths than the existing parametric bootstrap method proposed by Hall and Maiti (2006).
@@
When the number of area is small, there is a high likelihood to obtain negative estimates of the variance $A$ when applying Prasad Rao method. Throughout our simulation studies, we notice that the likelihood of obtaining negative estimates of variance might affect the performance of the parametric bootstrap methods in terms of both coverage probabilities and average lengths. One might consider a better variance estimator, such as the Fay-Herriot estimator we used, or a sensible truncation of negative estimates. In this study, we arbitrarily truncate the negative estimates of the variance $A$ at 0.01 as similar to Datta et al. (2005). To this end, in the future we will consider an extension of adjusted maximum likelihood estimators of $A$ such as the ones considered by Li and Lahiri (2010) or Hirose and Lahiri (2018) to the model proposed in this paper.